1. Two identical cars, mass 1.0 tonnes, traveling at 30.0 kph in opposite directions, collide. During the collision, the front of each the cars crumples and becomes 15 cm shorter. So let's say that each car decelerates to rest over 15 cm. What is the magnitude of the average acceleration?

a) Magnitude of average acceleration: _____ m.s^−2. ( Careful with significant figures. )
b) What is the magnitude of the average external force that produces this acceleration?
Magnitude of average external force: _____ kN.
1
Expert's answer
2020-03-11T14:13:00-0400

a)


"a=\frac{v^2}{2d}"

"a=\frac{(\frac{30}{3.6})^2}{2(0.15)}=231\frac{m}{s^2}"

b)


"F=ma=(1000)(231)=231\ kN"